# Selecting Critical Scenarios of DER Adoption in Distribution Grids Using Bayesian Optimization

### v1

@@ VI Conclusions and Future Work
In this work, we propose a new methodology to efficiently search through thousands of scenarios of PV adoption to identify those that lead to worst violations of bus voltages and line flows. Our Bayesian Optimization-driven search accurately navigates the high-dimensional scenario space, dramatically improving (2752×27-52$\times$27 - 52 × fewer evaluations) over a brute force approach that evaluates all simulated scenarios. We also demonstrate that the search is highly nontrivial and that greedy metrics like sorting by aggregate adopted PV capacity will miss most critical scenarios.
Several aspects of our methodology should be investigated further but are beyond the present scope. Further analysis of the best GP surrogates could offer faster initialization and re-fitting of the $f_{k}$’s; for example one could try to leverage the latent lower-dimensional structure hinted in Section V-B. It might be worthwhile to incorporate correlation among GP surrogates for related stress objectives. Another idea is to fit GPs directly for individual bus voltages $volt_{b}(\mathbf{x})$ in (2), rather than on groups of buses $\mathcal{B}_{k}$, and then group predictions. Some preliminary experiments showed promising gains in accuracy, although at a higher computational cost. Alternative acquisition functions might be warranted for very large feeders (over 1000 nodes) where $\alpha_{ND}$ becomes slow.

### v2

@@ VI Conclusions and Future Work
In this work, we propose a new methodology to efficiently search through thousands of scenarios of PV adoption to identify those that lead to worst violations of bus voltages and line flows. Our Bayesian Optimization-driven search accurately navigates the high-dimensional scenario space, dramatically improving ($27-52\times$ fewer evaluations) over a brute force approach that evaluates all simulated scenarios. We also demonstrate that the search is highly nontrivial and that naive heuristics like sorting by aggregate adopted PV capacity will miss most critical scenarios.
Several aspects of our methodology should be investigated further. Further analysis of the best GP surrogates could offer faster initialization and re-fitting of the $f_{k}$’s; for example one could try to leverage the latent lower-dimensional structure hinted in Section V-C. It might be worthwhile to incorporate correlation among GP surrogates to fuse learning of related stress objectives. Another idea is to fit GPs directly for individual bus voltages $volt_{b}(\mathbf{x})$ in (2), rather than on groups of buses $\mathcal{B}_{k}$, and then group predictions. Some preliminary experiments showed promising gains in accuracy, although at a higher computational cost. Alternative acquisition functions might be warranted for very large feeders (over 1000 nodes) where $\alpha_{ND}$ becomes slow.
The present work focused on distribution planning for radial grids. Further analysis is warranted for larger and/or meshed grids where more lines are likely to be critical. Finally, one could extend the methodology to consider multi-horizon planning where DER adoption evolves over multiple time steps and criticality is defined both locationally (what is violated) and temporally (when).